The perimeter of a triangle is 136. If the three sides of the triangle are (x+7), (2x-5), and (3x+2), what is the length of each side?
Answer by robertb(5830)

x+7 +2x-5 + 3x+2 = 136,
6x + 4 = 136,
6x = 132,
x = 22.
x + 7 = 29,
2x - 5 = 39,
3x + 2 = 68.